OpenStudy (anonymous):

If a cricket produces 30 chirps per minute at 50F and 58 chirps per minute at 62F, what is the slope of the graph of the linear equation describing the number of chirps per minute N as a function of the temperature T?

OpenStudy (anonymous):

find the slop of the line between (50,30)and (62,58) using \[m=\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (anonymous):

chirps per minute means you put chirps on top, minutes on the bottom

OpenStudy (anonymous):

\[\frac{58-30}{62-50}=\frac{28}{12}\] if i am not mistake. can reduce the fraction